# Asymmetric Perturbation in Solving Bilinear Saddle-Point Optimization

### v1

@@ Abstract
This paper proposes an asymmetric perturbation technique for solving saddle-point optimization problems, commonly arising in min-max problems, game theory, and constrained optimization. Perturbing payoffs or values are known to be effective in stabilizing learning dynamics and finding an exact solution or equilibrium. However, it requires careful adjustment of the perturbation magnitude; otherwise, learning dynamics converge to only an equilibrium. We establish an impossibility result that it almost never reaches an exact equilibrium as long as both players’ payoff functions are perturbed. To overcome this, we introduce an asymmetric perturbation approach, where only one player’s payoff function is perturbed. This ensures convergence to an equilibrium without requiring parameter adjustments, provided the perturbation strength parameter is sufficiently low. Furthermore, we empirically demonstrate fast convergence toward equilibria in both normal-form and extensive-form games.
@@ 1 Introduction
This paper proposes an asymmetric perturbation technique for solving saddle-point optimization problems, commonly arising in min-max problems, game theory, and constrained optimization. Over the past decade, no-regret learning algorithms have been extensively studied for computing (approximate) solutions or equilibria. When each player minimizes regret, the time-averaged strategies approximate Nash equilibria in two-player zero-sum games, that is, the algorithms guarantee average-iterate convergence. However, the actual sequence of strategies does not necessarily converge: it can cycle or diverge even in simple bilinear cases (Mertikopoulos et al., 2018; Bailey and Piliouras, 2018; Cheung and Piliouras, 2019). This is problematic, especially in large-scale games where neural networks are used for strategy representation, as averaging requires storing a separate model copy at each iteration.
@@
This is one of the main motivations for the study of last-iterate convergence, a stronger notion than average-iterate convergence, in which the strategies themselves converge to an equilibrium. One successful approach is to use optimistic learning algorithms, which essentially incorporate a one-step optimistic prediction that the environment will behave similarly in the next step. This idea has led to a number of effective algorithms, including Extra-Gradient methods (EG) (Liang and Stokes, 2019; Mokhtari et al., 2020), Optimistic Gradient Descent Ascent (OGDA) (Daskalakis and Panageas, 2019; Gidel et al., 2019; Mertikopoulos and Zhou, 2019), and Optimistic Multiplicative Weights Update (OMWU) (Daskalakis and Panageas, 2019; Lei et al., 2021a).
Alternatively, perturbing the payoffs by introducing strongly convex penalties (Facchinei and Pang, 2003) has long been recognized as a technique for achieving last-iterate convergence (Koshal et al., 2010; Tatarenko and Kamgarpour, 2019). This was typically done by decaying the perturbation magnitude over time, leading to asymptotic convergence. This idea has re-emerged through a different way of periodically updating the magnitude of perturbation, thereby stabilizing the learning dynamics (Perolat et al., 2021; Abe et al., 2022, 2023, 2024). Notably, such approaches have been successfully applied in practical settings, including learning in large-scale games (Perolat et al., 2022) and fine-tuning large language models via preference optimization (Munos et al., 2024), often in place of optimistic algorithms. However, these methods still require careful tuning of the perturbation strength; otherwise, the dynamics may converge only to an approximate equilibrium.
To clarify this limitation, we first prove an impossibility result showing that learning dynamics almost never reach an exact equilibrium when both players’ payoff functions are perturbed—a setting considered in most existing perturbation-based methods. In other words, symmetric regularization typically results in solutions that do not exactly match the equilibrium of the original game, regardless of how small the perturbation is.
@@
Motivated by this observation, we develop an asymmetric perturbation approach, where only one player’s payoff function is perturbed, while the other remains unperturbed. This simple modification yields a qualitatively different outcome: the learning dynamics converge to an exact saddle point, as long as the perturbation strength is sufficiently small—within a broad and practical range—without requiring fine-tuned parameter selection. Intuitively, leaving the opponent player $y$ unperturbed preserves the linearity of player $x$’s original objective, so adding a strongly convex perturbation does not significantly alter the solution. Consequently, even if strongly convex penalties are applied to the payoff function, the solution remains consistent between the original game and the perturbed game.
Furthermore, to demonstrate the effectiveness of our proposed approach, we provide two applications in normal-form and extensive-form games. First, we incorporate the technique into a gradient-based learning algorithm, which provably converges to a saddle point with a guaranteed rate. Empirical results on benchmark normal-form games show accelerated convergence. Second, we apply the technique to counterfactual regret (CFR) minimization (Zinkevich et al., 2007), a widely used method in extensive-form games, and demonstrate significant improvements in convergence speed on standard benchmarks.
@@ 2 Preliminaries
where $\mathcal{X}\subseteq\mathbb{R}^{m}$ (resp. $\mathcal{Y}\subseteq\mathbb{R}^{n}$) represents the $m$-dimensional (resp. $n$-dimensional) convex strategy space for player $x$ (resp. player $y$), and $A\in\mathbb{R}^{m\times n}$ is a game matrix. We assume that $\mathcal{X}$ and $\mathcal{Y}$ are polytopes. We refer to the function $x^{\top}Ay$ as the payoff function, and write $z=(x,y)$ as the strategy profile. This formulation includes many well-studied classes of games, such as two-player normal-form games and extensive-form games with perfect recall [^1] Two-player extensive-form games with perfect recall can be expressed as bilinear problems using sequence-form strategies (von Stengel, 1996).
@@
3 Asymmetric payoff perturbation
In this section, we explain our novel technique of asymmetric payoff perturbation. We first revisit the commonly used symmetric perturbation, where we show that the resulting solution of the perturbed game generally deviates from the equilibrium in the original game. Then, we demonstrate that a seemingly minor structural change—perturbing only one player’s payoff—can yield a dramatically different outcome: the solution of the perturbed game exactly matches the equilibrium in many cases.
3.1 Equilibrium shift under symmetric perturbation
@@
Let $x^{\mu}$ (resp. $y^{\mu}$) denote the minimax (resp. maximin) strategy in the symmetrically perturbed game (2) 222 The minimax strategy is uniquely determined because symmetrically perturbed games satisfy the strongly convex–strongly concave property., which can be solved at an exponentially fast rate (Cen et al., 2021, 2023; Pattathil et al., 2023; Sokota et al., 2023). It has also been shown that the distance between the solution of the symmetrically perturbed game $(x^{\mu},y^{\mu})$ and the solution in the original game $(x^{\ast},y^{\ast})$ is upper bounded by $\mathcal{O}(\mu)$ (Liu et al., 2023; Abe et al., 2024). However, they do not guarantee that the two solutions coincide, even for a small $\mu>0$. In contrast, the following theorem provides the first formal impossibility result, proving that $(x^{\mu},y^{\mu})$ almost never coincides with $(x^{\ast},y^{\ast})$.
Theorem 3.1.
The proof is provided in Appendix D. Additionally, we extend our analysis to the case where both players have different perturbation strengths, i.e., $\mu_{x}>0$ and $\mu_{y}>0$, as shown in Appendix B.
Discussion on Theorem 3.1.
(a) Symmetric Perturbation
(b) Asymmetric Perturbation
Figure 1: The proximity of $x^{\mu}$ to $x^{\ast}$ under the symmetric perturbation and the asymmetric perturbation with varying $\mu$. The game matrix $A$ is given by $[[0,1,-3],[-1,0,1],[3,-1,0]]$. The proximity to the minimax strategy is measured by the value of $\max_{y\in\mathcal{Y}}(x^{\mu})^{\top}Ay-v^{\ast}$.
When the game is symmetric.
Corollary 3.2.
Assume that $A^{\top}=-A$. Under the same setup as Theorem 3.1, the equilibrium $(x^{\mu},y^{\mu})$ in (2) always satisfies $x^{\mu}\neq x^{\ast}$ and $y^{\mu}\neq y^{\ast}$ for any $\mu>0$.
This is because it always holds that $v^{\ast}-\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)=0$ when $A^{\top}=-A$. Figure 1 shows the proximity of $x^{\mu}$ to $x^{\ast}$ with varying perturbation strength $\mu$ in a simple biased Rock-Paper-Scissors game. We observe that as long as $\mu>0$, $x^{\mu}$ remains distant from $x^{\ast}$. This observation supports the theoretical results in Theorem 3.1 and Corollary 3.2.
@@ 3.3 Equilibrium invariance under the asymmetric perturbation
In this section, we discuss the properties of the minimax strategies for asymmetrically perturbed games. Surprisingly, we can show that $x^{\mu}$ in (3) does correspond to a minimax strategy in the original game (1) for a sufficiently small $\mu$:
@@
Assume that the perturbation strength $\mu$ is set such that $\mu\in(0,\frac{\alpha}{\max_{x^{\ast}\in\mathcal{X}^{\ast}}\left\|x^{\ast}\right\|})$, where $\alpha>0$ is a constant depending only on $\mathcal{X}^{\ast}$. Then, the minimax strategy $x^{\mu}$ in the corresponding asymmetrically perturbed game (3) satisfies $x^{\mu}\in\mathcal{X}^{\ast}$.
Therefore, for all perturbation strengths $\mu$ less than a certain positive constant, $x^{\mu}$ coincides with the minimax strategy of the original game. Figure 1 visualizes this feature in a simple example. As long as the value of $\mu$ is less than $2.5$, the equilibrium strategy of the asymmetric perturbed game is invariant with respect to the original equilibrium strategy.
@@
The key element in proving Theorem 3.3 is the near-linear behavior of the objective function $g(x):=\max_{y\in\mathcal{Y}}x^{\top}Ay$ for player $x$ in the original game. Specifically, according to Wei et al. (2021), there exists a constant $\alpha>0$ such that:
$\forall x^{\ast}\in\mathcal{X}^{\ast},~{}g(x)-g\left(x^{\ast}\right)\geq\alpha\left\|x-\prod_{\mathcal{X}^{\ast}}(x)\right\|,$
@@
where $\prod_{\mathcal{A}}(a)=\mathop{\rm arg~{}min}\limits_{a^{\prime}\in\mathcal{A}}\left\|a-a^{\prime}\right\|$ represents the projection operator onto a given closed convex set $\mathcal{A}$. This inequality implies that deviating from the minimax strategy set $\mathcal{X}^{\ast}$ results in an increase in the objective function proportionally to the distance $\left\|x-\prod_{\mathcal{X}^{\ast}}(x)\right\|$. In contrast, the variation (i.e., the gradient) of the perturbation payoff function $\frac{\mu}{2}\left\|x\right\|^{2}$ can always be bounded by $\mathcal{O}(\mu)$ over $\mathcal{X}$. Hence, by choosing $\mu$ sufficiently small, we can ensure that the perturbation payoff function does not significantly incentivize player $x$ to deviate from $x^{\ast}$. In Figure 2, we illustrate this fact intuitively. The addition of the strongly convex function $\frac{\mu}{2}\|x\|^{2}$ does not shift the optimum $x^{*}$ of the original $g(x)$ if $\mu$ is sufficiently small, as the kink of the lines through $g(x)$ dominates. The detailed proof of Theorem 3.3 is provided in Appendix E.
@@
In summary, to compute a minimax strategy $x^{\ast}\in\mathcal{X}^{\ast}$ in the original game, it is sufficient to solve the asymmetrically perturbed game (3) with a small perturbation strength $\mu>0$. Note that, as mentioned above, we can also compute a maximin strategy $y^{\ast}\in\mathcal{Y}^{\ast}$ by solving the game $\max\limits_{y\in\mathcal{Y}}\min\limits_{x\in\mathcal{X}}\left\{x^{\top}Ay-\frac{\mu}{2}\left\|y\right\|^{2}\right\}$, where the payoff perturbation is applied only to player $y$.
@@ 4 Asymmetrically perturbed gradient descent ascent
This section proposes a gradient-based online learning algorithm, Asymmetrically Perturbed Gradient Descent Ascent (AsymP-GDA), for solving asymmetrically perturbed games (3). At each iteration $t\in[T]$, AsymP-GDA updates each player’s strategy according to the following update rule [^3] AsymP-GDA employs alternating updates rather than simultaneous updates, as recent work has demonstrated the advantages of the former over the latter (Lee et al., 2024).:
[…]
where $\eta>0$ is the learning rate. In AsymP-GDA, player $x$’s strategy $x^{t}$ is updated based on the gradient of the perturbed payoff function $x^{\top}Ay+\frac{\mu}{2}\left\|x\right\|^{2}$, while player $y$’s strategy $y^{t}$ is updated based on the gradient of the original payoff function $x^{\top}Ay$.
@@
Figure 3: Trajectories of strategies for player $x$ using AsymP-GDA, SymP-GDA, and GDA. The game matrix $A$ is set to $A=[[0,1,-3],[-1,0,1],[3,-1,0]]$, and the strategy spaces are set to $\mathcal{X}=\mathcal{Y}=\Delta^{3}$. The red point represents the minimax strategy in the original game. The trajectories originate from different initial strategies, demonstrating the learning dynamics under each method.

### v2

@@ Abstract
This paper proposes an asymmetric perturbation technique for solving bilinear saddle-point optimization problems, commonly arising in minimax problems, game theory, and constrained optimization. Perturbing payoffs or values is known to be effective in stabilizing learning dynamics and equilibrium computation. However, it requires decreasing perturbation magnitudes to ensure convergence to an equilibrium in the underlying game, resulting in a slower rate. To overcome this, we introduce an asymmetric perturbation approach, where only one player’s payoff function is perturbed. Exploiting the near-linear structure of bilinear problems, we show that, for a sufficiently small perturbation, the equilibrium strategy of the asymmetrically perturbed game coincides with an equilibrium strategy of the original game. Building on this property, we develop a perturbation-based learning algorithm with a linear last-iterate convergence rate to an equilibrium strategy of the original game, and we further show how to construct a parameter-free procedure that retains a linear rate. Finally, we empirically demonstrate fast convergence toward equilibria in both normal-form and extensive-form games.
@@ 1 Introduction
This paper proposes an asymmetric perturbation technique for solving saddle-point optimization problems, commonly arising in minimax problems, game theory, and constrained optimization. Over the past decade, no-regret learning algorithms have been extensively studied for computing (approximate) solutions or equilibria. When each player minimizes regret, the time-averaged strategies approximate Nash equilibria in two-player zero-sum games; that is, average-iterate convergence is guaranteed. However, the actual sequence of strategies does not necessarily converge and can cycle or diverge even in simple bilinear cases (Mertikopoulos et al., 2018; Bailey and Piliouras, 2018; Cheung and Piliouras, 2019). This is problematic, especially in large-scale games with neural network policies, since averaging requires storing a separate model at every iteration.
@@
This motivates the study of last-iterate convergence, a stronger notion than average-iterate convergence, in which the strategies themselves converge to an equilibrium. One successful approach is to use optimistic learning algorithms, which essentially incorporate a one-step optimistic prediction that the environment will behave similarly in the next step. This idea has led to several effective algorithms, including Extra-Gradient methods (EG) (Liang and Stokes, 2019; Mokhtari et al., 2020), Optimistic Gradient Descent Ascent (OGDA) (Daskalakis and Panageas, 2019; Gidel et al., 2019; Mertikopoulos and Zhou, 2019), and Optimistic Multiplicative Weights Update (OMWU) (Daskalakis and Panageas, 2019; Lei et al., 2021a). However, in large-scale settings where the gradient must be estimated from data or simulation, these algorithms can lose the last-iterate convergence property. For example, Abe et al. (2022) reports empirical non-convergence behavior under bandit feedback.
Alternatively, perturbing the payoffs with strongly convex penalties (Facchinei and Pang, 2003) has long been recognized as an effective technique for achieving last iterate convergence (Koshal et al., 2010; Tatarenko and Kamgarpour, 2019). This line of work has also shown strong performance in practical settings, including learning in large-scale games (Bakhtin et al., 2023) and fine-tuning large language models via preference optimization (Ye et al., 2024), often in place of optimistic algorithms. In prior work, the perturbation is almost always applied symmetrically, meaning both players’ payoff functions are perturbed by a strongly convex penalty. A known limitation is that, for any fixed perturbation strength $\mu>0$, the equilibrium of the symmetrically perturbed game generally remains only an approximation of an equilibrium of the original game, and the deviation scales with the strength of the perturbation (Liu et al., 2023; Abe et al., 2024). Consequently, recovering equilibria of the original game typically requires tuning $\mu$ carefully (e.g., as a function of the iteration budget) or employing a continuation-type procedure, which introduces a nontrivial trade-off between accuracy and convergence speed.
@@
To avoid these restrictions, we develop an asymmetric perturbation approach, in which only one player’s payoff function is perturbed while the other remains unperturbed. This simple modification yields a qualitatively different outcome. For any sufficiently small perturbation strength within a broad and practical range, the equilibrium strategy of the perturbed game coincides with that of the original game (see Theorem 3.1). Intuitively, leaving player $y$ unperturbed preserves the linearity of player $x$’s original objective, so adding a strongly convex perturbation does not significantly shift the solution (see Figure 2). Consequently, solving the asymmetrically perturbed game suffices to recover an equilibrium strategy of the original game.
Furthermore, to demonstrate the effectiveness of our findings, we first incorporate the asymmetric perturbation technique into a gradient-based learning algorithm for bilinear games, and show that it converges to a saddle point at a linear (exponentially fast) rate (see Theorem 4.1 and Corollary 4.2). This rate is provably faster than an $\tilde{\mathcal{O}}(1/t)$ [^1] We use $\tilde{\mathcal{O}}$ to denote a Landau notation that disregards a polylogarithmic factor. rate for existing symmetric perturbation-based approaches under the same bilinear setting (Liu et al., 2023). Moreover, leveraging the same invariance phenomenon, we show how to construct a parameter-free variant that retains a linear last-iterate convergence rate. Second, we apply our asymmetric perturbation to gradient-based learning methods for extensive-form games using a dilated regularizer (Hoda et al., 2010), a standard choice for learning over sequence-form strategy spaces (von Stengel, 1996). While our analysis focuses on bilinear games, the structural insight behind the asymmetric perturbation may extend beyond this setting, including two-player zero-sum Markov games, and serves as a bridge to the design of new perturbation-based learning algorithms.
@@ 2 Preliminaries
where $\mathcal{X}\subseteq\mathbb{R}^{m}$ (resp. $\mathcal{Y}\subseteq\mathbb{R}^{n}$) represents the $m$-dimensional (resp. $n$-dimensional) convex strategy space for player $x$ (resp. player $y$), and $A\in\mathbb{R}^{m\times n}$ is a game matrix. We assume that $\mathcal{X}$ and $\mathcal{Y}$ are polytopes. We refer to the function $x^{\top}Ay$ as the payoff function, and write $z:=(x,y)\in\mathcal{Z}:=\mathcal{X}\times\mathcal{Y}$ as the strategy profile. This formulation includes many well-studied classes of games, such as two-player normal-form games and extensive-form games with perfect recall.
@@
(a) Symmetric Perturbation
(b) Asymmetric Perturbation
Figure 1: The proximity of $x^{\mu}$ to $x^{\ast}$ under the symmetric perturbation and the asymmetric perturbation with varying $\mu$. The game matrix AA is given by [[0,1,3],[1,0,1],[3,1,0]][[0,1,-3],[-1,0,1],[3,-1,0]]. The proximity to the minimax strategy is measured by the value of $\max_{y\in\mathcal{Y}}(x^{\mu})^{\top}Ay-v^{\ast}$.
Symmetric perturbation.
@@
Let $x^{\mu}$ (resp. $y^{\mu}$) denote the minimax (resp. maximin) strategy in the symmetrically perturbed game (2)  [^2] The minimax strategy is uniquely determined because symmetrically perturbed games satisfy the strongly convex–strongly concave property., which can be solved at a linear rate (Cen et al., 2021, 2023; Pattathil et al., 2023; Sokota et al., 2023). It is known that the solution $(x^{\mu},y^{\mu})$ is only an approximation of an equilibrium of the original game, with an error bounded by $\mathcal{O}(\mu)$ (Liu et al., 2023; Abe et al., 2024). Consequently, typical perturbation-based methods must employ a decreasing schedule for $\mu$, or use an extremely small fixed $\mu$ tuned to the number of iterations $T$ (e.g., $\mu=\mathcal{O}(1/T)$), which requires careful hyperparameter tuning (Tatarenko and Kamgarpour, 2019; Bernasconi et al., 2022; Cai et al., 2023). See Figure 1 for a biased Rock–Paper–Scissors game where the perturbed solution differs from the original equilibrium. We provide a rigorous justification for this phenomenon in Appendix B, showing that for this instance, $(x^{\mu},y^{\mu})$ fails to coincide with any equilibrium of the original game for every fixed $\mu>0$.
3 Asymmetric Payoff Perturbation
In this section, we explain our novel technique of asymmetric payoff perturbation. We demonstrate that a seemingly minor structural change—perturbing only one player’s payoff—can yield a dramatically different outcome: the solution of the perturbed game exactly matches an equilibrium strategy $x^{\ast}$ in many cases.
@@ 3.2 Equilibrium Invariance under the Asymmetric Perturbation
In this section, we discuss the properties of the minimax strategies for asymmetrically perturbed games. Surprisingly, we can show that $x^{\mu}$ in (3) does correspond to a minimax strategy in the original game (1) for all $\mu$ smaller than a certain positive constant:
@@
Assume that the perturbation strength $\mu$ is set such that $\mu\in(0,\frac{\alpha}{\max_{x\in\mathcal{X}}\left\|x\right\|})$, where $\alpha>0$ is a constant depending only the game instance. Then, the minimax strategy $x^{\mu}$ in the corresponding asymmetrically perturbed game (3) satisfies $x^{\mu}\in\mathcal{X}^{\ast}$.
Thus, whenever $\mu$ is below a certain positive constant, $x^{\mu}$ coincides with the minimax strategy of the original game. Figure 1 illustrates this feature in a simple example (in that example, invariance holds for $\mu<2.5$). The asymmetric perturbation is closely related in spirit to Nesterov’s smoothing technique (Nesterov, 2005), which also perturbs only one side of a minimax formulation by adding a strongly convex term. A key distinction is which payoff is perturbed: smoothing is typically used by perturbing the opponent’s payoff on the maximization side, whereas we perturb the payoff of the player being optimized and leave the opponent’s payoff unchanged. This reversal is essential for Theorem 3.1, which identifies a regime of fixed $\mu$ in which the minimax strategy is invariant under the perturbation.
Remark 3.2 (Invariance without knowing the game-dependent constant).
The invariance result above is guaranteed when $\mu$ lies below a game-dependent upper bound, as characterized by $\alpha$. Nevertheless, this does not prevent us from exploiting invariance in practice for two reasons. First, we empirically observe that the invariance phenomenon appears to persist for a reasonably wide range of choices of $\mu$ in our experiments (see Figures 1 and 4). Second, in Section 4.2, we show how Theorem 3.1 can be leveraged to construct a parameter-free procedure that does not require knowledge of $\alpha$.
@@
The key ingredient in proving Theorem 3.1 is the near-linear behavior of the objective function $g(x):=\max_{y\in\mathcal{Y}}x^{\top}Ay$ for player $x$ in the original game. Specifically, according to Wei et al. (2021), there exists a constant $\alpha>0$ such that:
$\forall x^{\ast}\in\mathcal{X}^{\ast},~g(x)-g\left(x^{\ast}\right)\geq\alpha\left\|x-\Pi_{\mathcal{X}^{\ast}}(x)\right\|,$
@@
where $\Pi_{\mathcal{A}}(a)=\mathop{\rm arg~min}\limits_{a^{\prime}\in\mathcal{A}}\left\|a-a^{\prime}\right\|$ represents the projection operator onto a given closed convex set $\mathcal{A}$. This inequality implies that deviating from the minimax strategy set $\mathcal{X}^{\ast}$ results in an increase in the objective function proportionally to the distance $\left\|x-\Pi_{\mathcal{X}^{\ast}}(x)\right\|$. In contrast, the variation (i.e., the gradient) of the perturbation payoff function $\frac{\mu}{2}\left\|x\right\|^{2}$ can always be bounded by $\mathcal{O}(\mu)$ over $\mathcal{X}$. Hence, by choosing $\mu$ sufficiently small, we can ensure that the perturbation payoff function does not significantly incentivize player $x$ to deviate from $x^{\ast}$. In Figure 2, we illustrate this fact intuitively. The addition of the strongly convex function $\frac{\mu}{2}\|x\|^{2}$ does not shift the optimum $x^{*}$ of the original $g(x)$ if $\mu$ is sufficiently small, as the kink of the lines through $g(x)$ dominates. The detailed proof of Theorem 3.1 is provided in Appendix E.
@@
In summary, to compute a minimax strategy $x^{\ast}\in\mathcal{X}^{\ast}$ in the original game, it is sufficient to solve the asymmetrically perturbed game (3) with a small perturbation strength $\mu>0$. Note that, as mentioned above, one can also compute a maximin strategy $y^{\ast}\in\mathcal{Y}^{\ast}$ by solving the game $\max\limits_{y\in\mathcal{Y}}\min\limits_{x\in\mathcal{X}}\left\{x^{\top}Ay-\frac{\mu}{2}\left\|y\right\|^{2}\right\}$, where the payoff perturbation is applied only to player $y$.
Figure 3: Trajectories of strategies for player $x$ using AsymP-GDA, SymP-GDA, and GDA. The game matrix AA is set to A=[[0,1,3],[1,0,1],[3,1,0]]A=[[0,1,-3],[-1,0,1],[3,-1,0]], and the strategy spaces are set to $\mathcal{X}=\mathcal{Y}=\Delta^{3}$. The red point represents the minimax strategy in the original game. The trajectories originate from different initial strategies, demonstrating the learning dynamics under each method.
@@ 4 Asymmetrically Perturbed Gradient Descent Ascent
This section proposes a first-order method, Asymmetrically Perturbed Gradient Descent Ascent (AsymP-GDA), for solving asymmetrically perturbed games (3). At each iteration $t\in[T]$, AsymP-GDA updates each player’s strategy according to the following alternating updates [^3] AsymP-GDA employs alternating updates rather than simultaneous updates, as recent work has demonstrated the advantages of the former over the latter (Lee et al., 2024).:
$\begin{aligned} x^{t+1}&=\Pi_{\mathcal{X}}\left(x^{t}-\eta\left(Ay^{t}+\mu x^{t}\right)\right),\\ y^{t+1}&=\Pi_{\mathcal{Y}}\left(y^{t}+\eta A^{\top}x^{t+1}\right),\end{aligned}$ (6)
where $\eta>0$ is the learning rate. In AsymP-GDA, player $x$’s strategy $x^{t}$ is updated based on the gradient of the perturbed payoff function $x^{\top}Ay+\frac{\mu}{2}\left\|x\right\|^{2}$, while player $y$’s strategy $y^{t}$ is updated using the gradient of the original payoff function $x^{\top}Ay$. AsymP-GDA adds only negligible per-iteration runtime or memory overhead relative to standard alternating GDA, with the only additional operation being a single vector addition.